OpenStudy (anonymous):

Find the following cube root y^3 assume that letters can represent any real number

OpenStudy (anonymous):

that would be \[\sqrt[3]{y^3}=y\]

OpenStudy (anonymous):

nice use of "cube root" so you get a correct answer!

OpenStudy (anonymous):

this is like \[\sqrt[3]{27}=3\] because \[3^3=27\] you see that \[\sqrt[3]{y^3}=y\] because \[y^3=y^3\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

can I say 4 root of a number